150g of ice at 0°c is mixed with 300g of water at 50°c.calculate the temperature of the mixture.Take specific heat capacity of water as 4.2×10^3J/Kg/k (specific latent heat of fussion of ice as 3.34×10^5J/Kg/k
1
Expert's answer
2020-06-04T09:57:40-0400

The energy ballance for such situation will be the following:


"Q_{loss} = Q_{melt} + Q_{heat}"

where "Q_{loss}" is the energy lost by hot water, "Q_{melt}" is the energy required to melt the given amount of ice, "Q_{heat}" is the enrgy required to heat the cold water obtained after melting.

Writting the expression for "Q_{loss}"


"Q_{loss} = m_{hot}\cdot c_{water}(T_0 - T)"

where "m_{hot} = 0.3kg", "c_{water} = 4.2\cdot 10^{3} \dfrac{J}{kg\cdot K}", "T_0 = 50\degree C = 323K" and "T" is required final temperature.

Writting the expression for "Q_{melt}":


"Q_{melt} = m_{ice}\lambda"

where "m_{ice} = 0.15 g" and "\lambda = 3.34\cdot 10^5 \dfrac{J}{kg}".

The expression for "Q_{heat}" will be:


"Q_{heat} = m_{ice}\cdot c_{water}(T-273 K)"

Combining it all together:


"m_{hot}\cdot c_{water}(T_0 - T) = m_{ice}\lambda + m_{ice}\cdot c_{water}(T-273 K)"

Expressing "T" from the last equation, obtain:


"T = \dfrac{m_{hot}T_0 + m_{ice}\cdot 273K}{m_{hot} + m_{ice}} - \dfrac{m_{ice}\lambda}{c(m_{hot} + m_{ice})}"

Substituting numerical values:



"T = \dfrac{0.3\cdot 323 + 0.15\cdot 273}{0.3+0.15} - \dfrac{0.15\cdot 3.34\cdot 10^5}{ 4.2\cdot 10^{3}(0.3 + 0.15)} \approx 279.8K = 6.8 \degree C"

Answer. T = 6.8°C.